# FastVLM: Efficient Vision Encoding for Vision Language Models

### v1

@@ Abstract
Scaling the input image resolution is essential for enhancing the performance of Vision Language Models (VLMs), particularly in text-rich image understanding tasks. However, popular visual encoders such as ViTs become inefficient at high resolutions due to the large number of tokens and high encoding latency caused by stacked self-attention layers. At different operational resolutions, the vision encoder of a VLM can be optimized along two axes: reducing encoding latency and minimizing the number of visual tokens passed to the LLM, thereby lowering overall latency. Based on a comprehensive efficiency analysis of the interplay between image resolution, vision latency, token count, and LLM size, we introduce FastVLM—a model that achieves an optimized trade-off between latency, model size and accuracy. FastVLM incorporates FastViTHD, a novel hybrid vision encoder designed to output fewer tokens and significantly reduce encoding time for high-resolution images. Unlike previous methods, FastVLM achieves the optimal balance between visual token count and image resolution solely by scaling the input image, eliminating the need for additional token pruning and simplifying the model design. In the LLaVA-1.5 setup, FastVLM achieves 3.2$\times$ improvement in time-to-first-token (TTFT) while maintaining similar performance on VLM benchmarks compared to prior works. Compared to LLaVa-OneVision at the highest resolution (1152$\times$1152), FastVLM achieves comparable performance on key benchmarks like SeedBench and MMMU, using the same 0.5B LLM, but with 85$\times$ faster TTFT and a vision encoder that is 3.4$\times$ smaller.
@@ 1 Introduction
Vision Language Models (VLMs) enable visual understanding alongside textual inputs. VLMs are often built by passing visual tokens from a pretrained vision backbone to a pretrained Large Language Model (LLM) through a projection layer (also known as the connector module). Previous works [50, 49] have explored various training and fine-tuning strategies for these three components: the vision backbone, the adapter, and the LLM, which is typically a decoder-only model.
Several studies [56, 25, 60] highlight image resolution as a key factor in VLM performance, especially for text- and chart-rich data. However, increasing image resolution presents multiple challenges. First, pretrained vision encoders may not support high-resolution images, as this would make pretraining inefficient. To address this, one approach is to continuously pretrain the vision backbone to adapt it for high resolutions [6]. Alternatively, tiling strategies, such as Sphinx [48], S2 [66], and AnyRes [49], divide images into subregions, with each subregion processed independently by the backbone. This approach is particularly suitable for ViT-based backbones, which cannot accept varying input resolutions.
@@
In this work, we study VLM design and training from a runtime efficiency perspective motivated by their on-device deployment. We explore the optimization landscape as image resolution increases, aiming to improve accuracy-latency trade-off, where latency includes both the vision encoder inference time and the LLM prefilling time. Using extensive experiments with different LLM sizes and resolutions, we establish the Pareto optimal curve for a specific vision backbone, showing the best accuracy achievable within a given runtime budget (TTFT) based on different choices of resolution and LLM size.
We start by exploring the use of a hybrid convolutional-transformer architecture FastViT [76], pretrained with MobileCLIP [77], as a vision backbone for the VLM setup (Section 3.1). We demonstrate the potential of this hybrid backbone, which generates visual tokens over 4$\times$ faster than a ViT model while achieving higher overall VLM accuracy with multi-scale features (Section 3.1.1). However, further architectural optimization is possible when the primary goal is a high-resolution VLM (rather than embedding generation as in MobileCLIP-pretrained FastViT). We introduce a new hybrid vision encoder, FastViTHD, specifically designed for efficient VLM performance on high-resolution images (Section 3.2), and use it as the vision backbone to obtain FastVLM through visual instruction tuning. FastVLM demonstrates a significantly improved accuracy-latency trade-off over VLMs based on ViTs, convolutional encoders, and our previously discussed hybrid FastViT for different input image resolutions and LLM sizes (Figure 1(a), Figs. 1(b) and 4). In particular, FastVLM outperforms several prior works while being smaller, faster, and trained with less data (Table 6). Compared to LLaVa-OneVision [40] operating at the highest possible resolution (1152$\times$1152), FastVLM obtains comparable performance with the same 0.5B LLM, but with 85$\times$ faster TTFT and a 3.4$\times$ smaller vision encoder.
@@
  • We demonstrate the efficacy of hybrid vision backbones in VLMs compared to ViTs. We also introduce additional architectural interventions, such as multi-scale vision features, to further improve VLM performance while maintaining efficiency.
@@ 2 Related Works
Efficient Image Encoding. CLIP [63] pretrained vision transformers [22] are widely used for encoding images to a vision-language model. Some of the common choices include SigLIP [87], EVA-CLIP [69], InternViT [14] and DFN-CLIP [23]. To further boost performance, recent works like [32, 72, 67] use an ensemble of vision encoders pretrained using different objectives. These works are orthogonal to our work as they can benefit from using an efficient vision encoder among the ensemble of vision encoders. Since ViT architecture has been widely adopted in VLMs and the number of visual tokens remains one of the main causes of inefficiency, works like LLaVA-PruMerge [64] and Matryoshka-based token sampling methods [28, 7] have emerged to improve efficiency of image encoding by pruning visual tokens dynamically. Other works like [19, 9, 17, 18] reduce the number of visual tokens by incorporating perceiver-style resamplers or simply pooling using depth-wise convolutions or average pooling. Rather than using an isotropic architecture like ViT and then designing custom resamplers and projectors, hierarchical architectures can be a simpler design choice. Hierarchical backbones like ConvNeXT [53] and FastViT [76] produce fewer tokens as they downsample the input tensor at every stage of compute. Recently, ConvLLaVA [25] was introduced that uses a pure-convolutional vision encoder to encode images for a VLM. In our work, we introduce an improved convolution-transformer hybrid architecture for VLMs and discuss the pareto-optimal operating points when this architecture is scaled to higher input resolutions.
@@ 3 Architecture
Figure 2: Overview of the FastVLM architecture. FastVLM consists of our novel vision encoder, FastViTHD, trained using the same setup as LLaVa. The FastViTHD architecture is designed and trained for low latency at high resolution, utilizing novel multi-scale pooling, additional self-attention layers, and downsampling to generate 4$\times$ fewer tokens than FastViT, and 16$\times$ fewer tokens than ViT-L/14 at resolution 336.
In this section, we first explore the adoption of the FastViT hybrid vision encoder for vision-language modeling. We then introduce architectural interventions to improve performance on VLM tasks. We present FastViTHD, a new hybrid vision encoder designed for efficient high-resolution VLM. We provide comprehensive ablations to demonstrate the optimality of FastViTHD over FastViT and prior works for different LLMs and input resolutions. Figure 2 illustrates the overall architecture of FastVLM and FastViTHD. The training setup for all results in this section follows the same configuration as LLaVA-1.5 [49] with Vicuna-7B [90] as the LLM decoder, unless mentioned otherwise. See Sec. 4 for more details.
@@ 3.1 FastViT as VLM Image Encoder
We use a CLIP-pretrained hybrid vision encoder in our experiments. Specifically, we use the MCi2 image encoder from MobileCLIP [77], which has 35.7M parameters, is pretrained on DataCompDR [24, 77], and is based on the FastViT [76] architecture. For simplicity, we refer to this encoder as “FastViT” throughout the rest of the paper. As shown in Tab. 1, using FastViT at its CLIP-pretrained resolution (256$\times$256) alone does not yield a strong VLM. The main advantage of a hybrid encoder like FastViT lies in its favorable image resolution scaling characteristics, meaning it generates 5.2$\times$ fewer tokens than the ViT architecture with a patch size of 14. The considerable token reduction gives significant advantage to VLM, as it greatly reduces the prefilling time and time-to-first-token of the transformer decoders with quadratic-complexity. When the input resolution of FastViT is scaled to 768$\times$768, it produces the same number of visual tokens as ViT-L/14 with an input resolution of 336$\times$336 but achieves better performance on VLM benchmarks. This performance gap is even more pronounced on text-rich benchmarks like TextVQA and DocVQA, despite both architectures producing the same number of visual tokens. Moreover, even if it results in same number of tokens with higher resolution, it takes much less image encoding time, thanks to the efficient convolution layers.
@@
3.1.1 Multi-Scale Features
Typical convolutional and hybrid architectures split up the computations into 4 distinct stages with a downsampling operation between them. While the VLM relies on features from the penultimate layer, features in earlier stages of the network extract information at different granularity. Aggregating information from multiple scales can complement high-level features from the penultimate layer. This design is commonly used in object detection models like [47]. The architecture for multiple scale feature extraction is shown in Fig. 2. We ablate between 2 designs to pool features from different stages, i.e. AvgPooling and 2D Depthwise convolutions. From Tab. 2, we find that using depthwise convolutions results in better performance. Along with multi-scale features, we also experimented with different connector designs for FastViT (more details provided in supplementary materials). Collectively, these model interventions benefit hierarchical backbones like ConvNeXt and FastViT.
@@ 3.2 FastViTHD: High Resolution Encoder for VLM
While FastViT with the introduced model interventions performs well as an image encoder that is 8.7$\times$ smaller than ViT-L/14, previous studies [14, 42] have demonstrated that increasing the scale of the image encoder improves its generalization capabilities. Common practice in hybrid architectures is to scale the number of self-attention layers along with the width of each layer in a 4-stage architecture like [16, 85], but this approach has its drawbacks. From Fig. 3, simply scaling-up the number of self-attention layers in stage 3 and 4 by following [16, 85], in the existing FastViT architecture is not optimal. In fact, it is even slower than ConvNeXT-L. More details on the naively scaled version of FastViT are provided in supplementary materials. To reduce the impact of the added self-attention layers, we introduce an extra stage preceded by a downsampling layer, see Fig. 2. In this approach, the self-attention layers only handle tensors that have been downsampled by a factor of 32 on each side, compared to a factor of 16 in typical and more recent hybrid models like ViTamin [11]. The self-attention layers with the widest MLP layers process input tensors downsampled by a factor of 64 on each side. Our design reduces image encoding latency and generates 4$\times$ fewer tokens for the compute-intensive LLM decoder, thereby decreasing the time-to-first-token (TTFT). The architecture schematic is shown in Fig. 2, and we call this model FastViTHD.
@@ 3.2.1 Vision Encoder - Language Decoder Interplay
There are two approaches to input resolution scaling, the first approach is to change the input resolution of the model to the desired resolution. The second approach is to tile the input image and set the input resolution of the image encoder to the tile size. The tiled inference (AnyRes) was introduced in prior works [48, 50] to enable ViT models to process high resolution images. Since FastViTHD is designed to run inference efficiently on high input resolutions, we analyze the optimal operating point for various resolutions using the two strategies. From Fig. 6, we see that simply setting the input resolution of the model to the desired resolution results in VLMs with the best accuracy-latency tradeoff. Only at extremely high image resolutions like 1536$\times$1536 do we see the benefits of dynamic resolution, as model inference at this resolution is mostly affected by memory bandwidth available on-device. If dynamic resolution is desired, using a setting with fewer tiles exhibits better accuracy-latency tradeoff. With advancements in hardware and improvements in memory size and bandwidth, we expect that FastVLM can be efficiently scaled to even higher resolutions without the need for tiling strategies.
@@ 0.5B Model Comparison
R4 FastVLM (Ours) FastViTHD Qw.2 15+11.9 1024 256 125 166 62.9 80.8 61.6 87.0 63.6 31.0 78.7 66.2 31.7 68.9
@@ 1-2B Model Comparison
R5 MobileVLMv2 [18] ViT-L/14 ML. 1.2+3.6 336 144 304 458 59.3 66.7 52.1 84.3 - - - - - -
R6 FastVLM (Ours) FastViTHD Qw.2 15+1.1 768 144 125 152 63.9 75.8 64.4 87.2 65.2 35.4 79.4 61.3 34.9 71.7
R7 FastVLM (Ours) FastViTHD Qw.2 15+11.9 768 144 125 152 63.8 89.8 66.2 87.6 67.9 39.7 80.2 65.7 37.8 72.4
R8 DeepSeekVL [54] ViT-SO400M DS. - 384 576 430 - - - - 87.6 - 34.8 - - 32.2 66.7
R9 MM1 [60] ViT-H - 3000+1.5 1344 720 632 - - 62.3 68.2 87.4 67.5 39.4 - 68.4 33.2 65.6
R10 FastVLM (Ours) FastViTHD Qw.2 15+1.1 1024 256 125 233 64.2 74.8 66.0 88.0 66.9 37.6 79.9 67.7 33.1 71.4
R11 FastVLM (Ours) FastViTHD Qw.2 15+11.9 1024 256 125 233 63.9 90.5 68.4 87.6 72.9 41.0 80.8 72.5 38.4 73.0
@@
R12 InstructBLIP [19] ViT-g/14 Vic. 129+1.2 224 32 1012 302 49.2 60.5 50.1 - 60.9 26.2 - - 30.6 -
R13 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 256 16 125 150 60.6 69.2 53.1 82.3 60.4 27.5 74.7 17.4 36.2 63.7
R14 FastVLM (Ours) FastViTHD Vic. 15+1.1 256 16 125 150 62.1 75.7 57.2 83.9 64.0 31.5 77.3 29.8 37.6 68.8
R15 MobileVLMv2 [18] ViT-L/14 Vic. 1.2+3.6 336 144 304 460 62.6 74.8 62.3 85.3 - - - - - -
R16 ConvLLaVA [25] ConvNeXT-L Vic. 4.9+0.6 768 144 200 496 - - 59.1 87.3 - 44.8 - 44.8 36.3 68.8
R17 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 768 144 125 387 62.4 67.6 62.9 87.7 63.8 31.5 78.9 32.9 34.9 68.2
R18 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 768 144 125 387 63.2 73.5 67.5 86.3 63.9 33.0 79.1 57.3 36.9 69.9
R19 FastVLM (Ours) FastViTHD Vic. 15+1.1 768 144 125 387 65.0 78.7 69.4 87.5 67.0 42.2 81.3 65.5 37.0 73.7
R20 FastVLM (Ours) FastViTHD Qw.2 15+1.1 768 144 125 446 65.6 85.9 69.5 87.2 73.0 41.3 81.3 66.9 43.6 75.3
R21 FastVLM (Ours) FastViTHD Qw.2 15+11.9 768 144 125 446 64.5 96.2 72.8 87.8 73.9 45.4 81.9 72.0 48.1 74.7
R22 Qwen-VL [4] ViT-G/14 Qw. 1400+50 448 256 1844 - 59.3 67.1 63.8 - - - 79.5 65.1 - -
R23 Qwen-VL-Chat [4] ViT-G/14 Qw. 1400+50 448 256 1844 - 57.5 68.2 61.5 - - - 78.2 62.6 - -
R24 ConvLLaVA [25] ConvNeXT-L Vic. 4.9+0.6 1024 256 200 1157 - - 62.5 87.7 - 44.4 - 48.5 35.1 69.3
R25 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 1024 256 125 577 63.1 67.4 64.4 88.1 64.8 31.7 79.2 35.6 35.1 68.5
R26 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 1024 256 125 577 63.3 74.1 67.4 87.1 66.5 32.4 79.3 62.8 37.3 69.9
R27 FastVLM (Ours) FastViTHD Vic. 15+1.1 1024 256 125 577 65.2 80.3 70.6 87.2 71.5 40.1 81.6 72.4 36.7 73.5
R28 FastVLM (Ours) FastViTHD Qw.2 15+1.1 1024 256 125 641 65.8 84.9 72.1 87.8 75.8 44.1 81.7 73.3 46.2 75.1
R29 LLaVA-1.5 [49] ViT-L/14 Vic. 0.5+0.6 336 576 304 1297 62.0 70.4 58.2 85.9 59.6 31.1 76.6 28.1 35.3 66.1
R30 MobileVLMv2 [18] ViT-L/14 Vic. 1.2+3.6 336 576 304 1297 64.6 74.8 66.8 86.1 - - - - - -
R31 ShareGPT4V [12] ViT-L/14 Vic. 1.2+0.7 336 576 304 1297 63.3 68.4 60.4 85.7 72.6 37.6 80.6 - - 69.7
R32 ViTamin [11] ViTamin-L Vic. 0.5+0.6 384 576 333 1308 61.6 67.6 59.8 85.5 66.1 33.6 78.9 - - -
R33 ConvLLaVA [25] ConvNeXT-L Vic. 4.9+0.6 1536 576 200 2740 - - 65.8 87.3 - 45.9 - 59.0 35.8 70.2
R34 VILA [45] ViT-L/14 L-2 50+1 336 576 304 1297 62.3 68.2 64.4 85.5 69.7 34.9 79.9 - - 62.8
R35 LLaVA-FlexAttn [41] ViT-L/14 Vic. 0.5+0.6 1008 576 304 - 62.2 - 48.9 85.9 - 29.4 78.7 - - -
R36 MM1 [60] ViT-H - 3000+1.5 1344 720 632 - - 72.6 72.8 86.6 81.5 42.1 82.8 76.8 37.0 69.9
R37 LLaVA-NeXT ViT-L/14 L-3 - 672 2880 304 20347 65.2 72.8 64.6 - 80.1 - - 78.2 41.7 72.7
R38 FastVLM (Ours) FastViTHD Qw.2 15+6.5 1024 256 125 641 66.0 87.4 73.1 87.3 72.4 47.6 82.3 78.7 42.8 75.9
R39 FastVLM (Ours) FastViTHD Qw.2 15+11.9 1024 256 125 641 64.5 96.0 74.8 87.2 73.5 49.9 82.3 78.9 46.4 75.1

### v2

@@ Abstract
Scaling the input image resolution is essential for enhancing the performance of Vision Language Models (VLMs), particularly in text-rich image understanding tasks. However, popular visual encoders such as ViTs become inefficient at high resolutions due to the large number of tokens and high encoding latency. At different operational resolutions, the vision encoder of a VLM can be optimized along two axes: reducing encoding latency and minimizing the number of visual tokens passed to the LLM, thereby lowering overall latency. Based on a comprehensive efficiency analysis of the interplay between image resolution, vision latency, token count, and LLM size, we introduce FastVLM—a model that achieves an optimized trade-off between resolution, latency, and accuracy. FastVLM incorporates FastViTHD, a novel hybrid vision encoder designed to output fewer tokens and significantly reduce encoding time for high-resolution images. Unlike previous methods, FastVLM achieves the optimal balance between visual token count and image resolution solely by scaling the input image, eliminating the need for additional token pruning and simplifying the model design. In the LLaVA-1.5 setup, FastVLM achieves 3.2$\times$ improvement in time-to-first-token (TTFT) while maintaining similar performance on VLM benchmarks compared to prior works. Compared to LLaVa-OneVision at the highest resolution (1152$\times$1152), FastVLM achieves better performance on key benchmarks like SeedBench, MMMU and DocVQA, using the same 0.5B LLM, but with 85$\times$ faster TTFT and a vision encoder that is 3.4$\times$ smaller. Code and models are available at https://github.com/apple/ml-fastvlm.
@@ 1 Introduction
Vision Language Models (VLMs) enable visual understanding alongside textual inputs. VLMs are often built by passing visual tokens from a pretrained vision backbone to a pretrained Large Language Model (LLM) through a projection layer. Previous works [54, 53] have explored various training and fine-tuning strategies for these three components: the vision backbone, the projection layer, and the LLM, which is typically a decoder-only transformer [84] model.
Several studies [61, 28, 66] highlight image resolution as a key factor in VLM performance, especially for text and chart-rich data. However, increasing image resolution presents multiple challenges. First, pretrained vision encoders may not support high-resolution images, as this would make pretraining inefficient. To address this, one approach is to continuously pretrain the vision backbone to adapt it for high resolutions [6]. Alternatively, tiling strategies, such as Sphinx [52], S2 [72], and AnyRes [53], divide images into subregions, with each subregion processed independently by the backbone.
@@
In this work, we study VLM design and training from a runtime efficiency perspective. We explore the optimization landscape as image resolution increases, aiming to improve accuracy-latency trade-off, where latency includes both the vision encoder inference time and the LLM prefilling time. Using extensive experiments with different LLM sizes and resolutions, we establish the Pareto optimal curve for a specific vision backbone, showing the best accuracy achievable within a given runtime budget (TTFT) based on different choices of resolution and LLM size.
We start by exploring the use of a hybrid convolutional-transformer architecture FastViT [82], pretrained with MobileCLIP [83], as a vision backbone for the VLM setup (Section 3.1). We demonstrate the potential of this hybrid backbone, which generates visual tokens over 4$\times$ faster than a ViT model while achieving higher overall VLM accuracy with multi-scale features (Section 3.1.1).
However, further architectural optimization is possible when the primary goal is a high-resolution VLM (rather than embedding generation as in MobileCLIP-pretrained FastViT). We introduce a new hybrid vision encoder, FastViTHD, specifically designed for efficient VLM performance on high-resolution images (Section 3.2), and use it as the vision backbone to obtain FastVLM through visual instruction tuning. FastVLM demonstrates a significantly improved accuracy-latency trade-off over VLMs based on ViTs, convolutional encoders, and our previously discussed hybrid FastViT for different input image resolutions and LLM sizes (Figure 1(a), Figs. 1(b) and 4). In particular, FastVLM outperforms several prior works while being smaller, faster, and trained with less data (Table 6). Compared to LLaVA-OneVision [45] operating at the highest possible resolution (1152$\times$1152), FastVLM obtains comparable performance with the same 0.5B LLM, but with 85$\times$ faster TTFT and a 3.4$\times$ smaller vision encoder.
@@
  • We show that hybrid vision backbones outperform ViTs in VLMs, and introduce additional architectural interventions, such as multi-scale vision features, to further improve VLM performance while maintaining efficiency.
@@ 2 Related Works
Efficient Image Encoding. CLIP [69] pretrained vision transformers [24] are widely used for encoding images in vision-language models, with popular choices including SigLIP [94], EVA-CLIP [75], InternViT [15] and DFN-CLIP [26]. To enhance performance, recent works [36, 78, 73] employ ensembles of vision encoders trained with different objectives. These works are orthogonal to our work as they can benefit from using an efficient vision encoder among the ensemble of vision encoders. Since ViT-based architectures are a popular choice for VLMs, inefficiencies arise from the number of visual tokens outputted by the encoder, prompting methods like LLaVA-PruMerge [70] and Matryoshka-based token sampling [32, 7] to dynamically prune tokens. Other approaches [20, 9, 18, 19] reduce tokens using perceiver-style resamplers or pooling techniques. Rather than using an isotropic architecture like ViT and then designing custom resamplers and projectors, hierarchical architectures can be a simpler design choice. Hierarchical backbones like ConvNeXT [57] and FastViT [82] produce fewer tokens as they downsample the input tensor at every stage of compute. Recently, ConvLLaVA [28] was introduced that uses a pure-convolutional vision encoder to encode images for a VLM. In our work, we introduce an improved convolution-transformer hybrid architecture for VLMs and discuss the Pareto optimal operating points when this architecture is scaled to higher input resolutions.
@@ 3 Architecture
Figure 2: Overview of the FastVLM architecture. FastVLM consists of our novel vision encoder, FastViTHD, trained using the same setup as LLaVA. The FastViTHD architecture is designed for low latency at high resolution, by utilizing additional self-attention layers, and downsampling to generate 4$\times$ fewer tokens than FastViT, and 16$\times$ fewer tokens than ViT-L/14 at resolution 336.
In this section, we first explore the adoption of the FastViT hybrid vision encoder for vision-language modeling. We then introduce architectural interventions to improve performance on VLM tasks. We present FastViTHD, a new hybrid vision encoder designed for an efficient high-resolution VLM. We provide comprehensive ablations to demonstrate the optimality of FastViTHD over FastViT and prior works for different LLMs and input resolutions. Figure 2 illustrates the overall architecture of FastVLM and FastViTHD. The training setup for all results in this section follows the same configuration as LLaVA-1.5 [53] with Vicuna-7B [98] as the LLM decoder, unless mentioned otherwise. See Sec. 4 for more details.
@@ 3.1 FastViT as VLM Image Encoder
We use a CLIP-pretrained hybrid vision encoder, specifically the MCi2 image encoder from MobileCLIP [83], which has 35.7M parameters and is based on the FastViT architecture. For simplicity, we refer to this encoder as “FastViT” throughout the rest of the paper. As shown in Tab. 1, using FastViT at its CLIP-pretrained resolution (256$\times$256) alone does not yield a strong VLM. The main advantage of a hybrid encoder like FastViT lies in its favorable image resolution scaling characteristics, meaning it generates 5.2$\times$ fewer tokens than the ViT architecture with a patch size of 14. The token reduction gives significant advantage to VLMs, as it reduces the prefilling time and time-to-first-token (TTFT) of the transformer decoders. When the input resolution of FastViT is scaled to 768$\times$768, it produces the same number of visual tokens as ViT-L/14 with an input resolution of 336$\times$336 but achieves better performance on VLM benchmarks. This performance gap is even more pronounced on text-rich benchmarks like TextVQA and DocVQA, despite both architectures producing the same number of visual tokens. Moreover, even with the same token count at higher resolution, it encodes images much faster due to efficient convolution layers.
@@
3.1.1 Multi-Scale Features
Typical convolutional and hybrid architectures split up the computations into 4 distinct stages with a downsampling operation between them. While the VLM relies on features from the penultimate layer, features in earlier stages of the network extract information at different granularity. Aggregating information from multiple scales can complement high-level features from the penultimate layer.
The architecture for multiple scale feature extraction is shown in Fig. 2. We ablate between 2 designs to pool features from different stages, i.e. AvgPooling and 2D Depthwise convolutions. From Tab. 2, we find that using depthwise convolutions results in better performance.
@@ 3.2 FastViTHD: High Resolution Encoder for VLM
While FastViT with the introduced model interventions performs well as an image encoder that is 8.7$\times$ smaller than ViT-L/14, previous studies [15, 47] have demonstrated that increasing the scale of the image encoder improves its generalization capabilities.
Hybrid architectures [17, 92] typically scale the number of self-attention layers and width in a 4-stage design, but this has drawbacks. From Fig. 3, simply scaling-up the number of self-attention layers in stages 3 and 4 of FastViT, as done in prior works, is suboptimal and slower than ConvNeXT-L. To mitigate this, we introduce an extra stage with a downsampling layer, ensuring self-attention operates on tensors downsampled by a factor of 32, rather than 16 as in recent models like ViTamin, see Fig. 2. More details on the naive scaling approach can be found in Sec.B. Our design reduces image encoding latency and generates 4$\times$ fewer tokens for the compute-intensive LLM decoder, thereby decreasing the time-to-first-token (TTFT). The architecture schematic is shown in Fig. 2, and we call this model FastViTHD.
@@ 3.2.1 Vision Encoder - Language Decoder Interplay
There are two ways to scale input resolution: adjusting the model’s input resolution directly or tiling the image and setting the encoder’s resolution to the tile size. The tiled inference (AnyRes) was introduced in prior works [52, 54] to enable ViT models to process high resolution images. Since FastViTHD is designed to run inference efficiently on high input resolutions, we analyze the optimal operating point for various resolutions using the two strategies. From Fig. 6, we see that setting the model’s input resolution directly to the desired resolution offers the best accuracy-latency tradeoff, with dynamic resolution benefiting only at extreme resolutions like 1536$\times$1536, due to memory bandwidth limitations. If dynamic resolution is desired, using a setting with fewer tiles exhibits better accuracy-latency tradeoff. Further discussion on this setup is presented in Sec. C.1.
@@ 0.5B Model Comparison
R4 FastVLM (Ours) FastViTHD Qw.2 15+12.5 1024 256 125 166 63.1 81.5 62.9 86.6 66.7 29.8 78.8 70.4 32.9 69.2
R5 FastVLM (Ours) FastViTHD Qw.2 15+23.1 1024 256 125 166 62.7 82.0 65.8 86.2 - 37.5 78.6 79.1 33.6 69.3
@@ 1-2B Model Comparison
R6 MobileVLMv2 [19] ViT-L/14 ML. 1.2+3.6 336 144 304 458 59.3 66.7 52.1 84.3 - - - - - -
R7 FastVLM (Ours) FastViTHD Qw.2 15+1.1 768 144 125 152 63.9 75.8 64.4 87.2 65.2 35.4 79.4 61.3 34.9 71.7
R8 FastVLM (Ours) FastViTHD Qw.2 15+12.5 768 144 125 152 64.2 87.9 66.4 88.3 68.5 41.1 80.1 69.4 38.1 72.4
R9 DeepSeekVL [58] ViT-SO400M DS. - 384 576 430 - - - - 87.6 - 34.8 - - 32.2 66.7
R10 MM1 [66] ViT-H - 3000+1.5 1344 720 632 - - 62.3 68.2 87.4 67.5 39.4 - 68.4 33.2 65.6
R11 FastVLM (Ours) FastViTHD Qw.2 15+1.1 1024 256 125 233 64.2 74.8 66.0 88.0 66.9 37.6 79.9 67.7 33.1 71.4
R12 FastVLM (Ours) FastViTHD Qw.2 15+12.5 1024 256 125 233 64.3 89.9 69.0 87.8 68.0 43.6 80.7 75.6 39.1 73.1
R13 FastVLM (Ours) FastViTHD Qw.2 15+23.1 1024 256 125 233 64.3 92.9 71.2 87.8 - 51.0 80.9 84.2 38.0 72.6
@@
R14 InstructBLIP [20] ViT-g/14 Vic. 129+1.2 224 32 1012 302 49.2 60.5 50.1 - 60.9 26.2 - - 30.6 -
R15 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 256 16 125 150 60.6 69.2 53.1 82.3 60.4 27.5 74.7 17.4 36.2 63.7
R16 FastVLM (Ours) FastViTHD Vic. 15+1.1 256 16 125 150 62.1 75.7 57.2 83.9 64.0 31.5 77.3 29.8 37.6 68.8
R17 MobileVLMv2 [19] ViT-L/14 Vic. 1.2+3.6 336 144 304 460 62.6 74.8 62.3 85.3 - - - - - -
R18 ConvLLaVA [28] ConvNeXT-L Vic. 4.9+0.6 768 144 200 496 - - 59.1 87.3 - 44.8 - 44.8 36.3 68.8
R19 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 768 144 125 387 62.4 67.6 62.9 87.7 63.8 31.5 78.9 32.9 34.9 68.2
R20 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 768 144 125 387 63.2 73.5 67.5 86.3 63.9 33.0 79.1 57.3 36.9 69.9
R21 FastVLM (Ours) FastViTHD Vic. 15+1.1 768 144 125 387 65.0 78.7 69.4 87.5 67.0 42.2 81.3 65.5 37.0 73.7
R22 FastVLM (Ours) FastViTHD Qw.2 15+1.1 768 144 125 446 65.6 85.9 69.5 87.2 73.0 41.3 81.3 66.9 43.6 75.3
R23 FastVLM (Ours) FastViTHD Qw.2 15+12.5 768 144 125 446 64.7 96.0 71.9 87.4 73.6 49.4 81.1 75.3 44.6 74.3
R24 Qwen-VL [4] ViT-G/14 Qw. 1400+50 448 256 1844 - 59.3 67.1 63.8 - - - 79.5 65.1 - -
R25 Qwen-VL-Chat [4] ViT-G/14 Qw. 1400+50 448 256 1844 - 57.5 68.2 61.5 - - - 78.2 62.6 - -
R26 ConvLLaVA [28] ConvNeXT-L Vic. 4.9+0.6 1024 256 200 1157 - - 62.5 87.7 - 44.4 - 48.5 35.1 69.3
R27 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 1024 256 125 577 63.1 67.4 64.4 88.1 64.8 31.7 79.2 35.6 35.1 68.5
R28 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 1024 256 125 577 63.3 74.1 67.4 87.1 66.5 32.4 79.3 62.8 37.3 69.9
R29 FastVLM (Ours) FastViTHD Vic. 15+1.1 1024 256 125 577 65.2 80.3 70.6 87.2 71.5 40.1 81.6 72.4 36.7 73.5
R30 FastVLM (Ours) FastViTHD Qw.2 15+1.1 1024 256 125 641 65.8 84.9 72.1 87.8 75.8 44.1 81.7 73.3 46.2 75.1
R31 LLaVA-1.5 [53] ViT-L/14 Vic. 0.5+0.6 336 576 304 1297 62.0 70.4 58.2 85.9 59.6 31.1 76.6 28.1 35.3 66.1
R32 MobileVLMv2 [19] ViT-L/14 Vic. 1.2+3.6 336 576 304 1297 64.6 74.8 66.8 86.1 - - - - - -
R33 ShareGPT4V [13] ViT-L/14 Vic. 1.2+0.7 336 576 304 1297 63.3 68.4 60.4 85.7 72.6 37.6 80.6 - - 69.7
R34 ViTamin [12] ViTamin-L Vic. 0.5+0.6 384 576 333 1308 61.6 67.6 59.8 85.5 66.1 33.6 78.9 - - -
R35 ConvLLaVA [28] ConvNeXT-L Vic. 4.9+0.6 1536 576 200 2740 - - 65.8 87.3 - 45.9 - 59.0 35.8 70.2
R36 VILA [50] ViT-L/14 L-2 50+1 336 576 304 1297 62.3 68.2 64.4 85.5 69.7 34.9 79.9 - - 62.8
R37 LLaVA-FlexAttn [46] ViT-L/14 Vic. 0.5+0.6 1008 576 304 - 62.2 - 48.9 85.9 - 29.4 78.7 - - -
R38 MM1 [66] ViT-H - 3000+1.5 1344 720 632 - - 72.6 72.8 86.6 81.5 42.1 82.8 76.8 37.0 69.9
R39 LLaVA-NeXT ViT-L/14 L-3 - 672 2880 304 20347 65.2 72.8 64.6 - 80.1 - - 78.2 41.7 72.7
R40 FastVLM (Ours) FastViTHD Qw.2 15+6.5 1024 256 125 641 66.0 87.4 73.1 87.3 72.4 47.6 82.3 78.7 42.8 75.9
R41 FastVLM (Ours) FastViTHD Qw.2 15+12.5 1024 256 125 641 65.2 95.7 73.4 86.9 71.1 48.4 81.6 82.7 47.3 74.1
R42 FastVLM (Ours) FastViTHD Qw.2 15+23.1 1024 256 125 641 64.1 96.8 76.1 87.2 - 56.3 81.9 88.0 44.0 74.5